what is the half life of a compound that decays at the formula y=ae^-0.0736t
Answer by jim_thompson5910(35256)

Plug in y = a/2 and solve for t


y = a*e^(-0.0736t)

a/2 = a*e^(-0.0736t)

1/2 = e^(-0.0736t)

ln(1/2) = ln(e^(-0.0736t))

ln(1/2) = -0.0736t*ln(e)

ln(1/2) = -0.0736t*1

ln(1/2) = -0.0736t

ln(1/2)/(-0.0736) = t

9.41776060543403 = t

t = 9.41776060543403


The halflife is approximately 9.41776060543403 (use whatever units are part of the problem)